Integrate the following function w.r.t. x and find within the limit x=2 to x=4:
a) X3
b)(10+ x2)
1
Expert's answer
2020-07-29T09:43:14-0400


a)


"\int{x^3}dx=(\frac{x^4}{4})_2^4=\frac{4^4}{4}-\frac{2^4}{4}=64-4=60"


b)


"\int({10+x^2})dx=(10x+\frac{x^3}{3})_2^4="


"=10\cdot 4+\frac{4^3}{3}-10\cdot 2-\frac{2^3}{3}=38.67"